# Mapless Collision-Free Flight via MPC using Dual KD-Trees in Cluttered Environments

### v2

@@ I Introduction
While effective in real-world cluttered environments, this cascade paradigm faces two inherent limitations. First, the accumulated error from each of the sequential modules including occupancy mapping, path planning, trajectory generation, and trajectory tracking could produce biased actions making the quadrotor clash[7]. Second, the sequential workflow also introduces significant computational delays, impairing responsiveness in high-speed scenarios[8]. Both limitations severely restrict the agility and safety of quadrotors. It naturally raises the question of whether approaches without mapping or trajectory generation exist such that agility and safety can be significantly enhanced.
@@
Specifically, we introduce a dual KD-tree [15, 16, 17, 18, 19] mechanism. The Obstacle KD-Tree enables real-time nearest-neighbor queries for collision cost calculation, while the Edge KD-Tree optimizes waypoint adjustment by integrating dilated obstacle boundaries, providing robust initial guesses for the MPC solver to improve optimization efficiency and avoid local minima. This dual-KD tree structure not only simplifies the computational pipeline but also enhances the framework’s ability to handle complex environments, ensuring both dynamic feasibility and real-time performance. Our method eliminates the need for explicit trajectory generation, tracking, or complex mapping processes, significantly enhancing the agility and robustness of autonomous flight in cluttered environments.
@@ III System Overview
The system begins by sparsely sampling a fixed number of waypoints from the current position toward the goal. It then constructs the Obstacle KD-Tree and Edge KD-Tree using the depth map from the RGB-D camera and adjusts the waypoints if a potential collision is detected. The Obstacle KD-Tree stores obstacle points extracted from the depth map, while the Edge KD-Tree is built from a dilated, edge-filtered version of the depth map. Next, the MPC optimization module refines the waypoints generated by the Edge KD-Tree, incorporating kinematic constraints and collision costs. The optimization process minimizes a cost function over a fixed finite time horizon, producing an optimal sequence of control inputs and predictive waypoints. The resulting control commands are then sent to the quadrotor’s low-level flight controller (FCU). This entire process repeats at each time step.
@@ IV Dual KD-Tree Perception Module
Below, we provide a detailed description of the construction and query process of the dual KD-Tree module. The input depth map is first downsampled to a resolution of $64\times 48$ pixels to reduce computational load. Two spatially indexed structures are then constructed to enable efficient collision checking and trajectory refinement:
@@ IV-1 Edge KD-Tree and Waypoint Coarse Adjustment
[…]
@@ V-A Optimization Formulation
s.t. $x_{h+1}=\text{RK4}\left(f_{d}\left(x_{h},u_{h},\theta\right),\Delta t\right)$
$u_{\text{min}}\leq u_{h}\leq u_{\text{max}}$
$x_{1}=x_{\text{init}}$
[…]
@@ V-B Cost Function
The cost function $\mathcal{C}\left(x_{h},u_{h}\right)$ is a weighted sum of four terms: collision avoidance cost $c_{a}$, terminal cost $c_{goal}$, waypoint following cost $c_{wp}$, and smoothness cost $c_{u}$. The cost function is defined as:
@@ V-B1 Collision Avoidance Cost
The collision avoidance cost $c_{a}(x_{h})$ penalizes the proximity of the h-th waypoint to obstacles. We compute this cost using distances between the quadrotor position $p_{h}$ and $M$ nearest obstacle points $\{o_{h}^{i}\}_{i=1}^{M}$ queried from the Obstacle KD-Tree:
$c_{a}(x_{h})$ $=\lambda_{c}\sum_{i=1}^{M}\lVert v_{h}^{i}\rVert\ln\left[1+e^{\beta(r_{q}-\lVert p_{h}-o_{h}^{i}\rVert)}\right]$ (3)
@@ V-B2 Terminal Cost
The terminal cost $\mathcal{C}_{goal}\left(x_{H},u_{H}\right)$ is used to guide the quadrotor to the final goal point. This cost is computed with the goal state $x_{\text{goal}}$:
$\mathcal{C}_{goal}\left(x_{H}\right)=\left(x_{H}-x_{\text{goal}}\right)^{T}Q_{\text{goal}}\left(x_{H}-x_{\text{goal}}\right)$ (4)
Here, $Q_{\text{goal}}$ is a diagonal weighting matrix that adjusts the relative importance of the terminal cost. And the goal state $x_{\text{goal}}$ is the final waypoint in the sequence.
@@ V-B3 Discrete Waypoint Following Cost
The discrete waypoint following cost, $c_{wp}(x_{h})$, encourages the discretized states $x_{h}$ to closely follow the sparse waypoints $x_{r}^{h}$:
$c_{wp}(x_{h})=\left(x_{h}-x_{r}^{h}\right)^{T}Q_{wp}\left(x_{h}-x_{r}^{h}\right)$ (5)
Here, $Q_{wp}$ is a diagonal weighting matrix. Adjusting $Q_{wp}$ influences how tightly the quadrotor follows the waypoints.
@@ V-B4 Smoothness Cost
The smoothness cost $c_{u}(u_{h})$ penalizes large control inputs $u_{h}$, promoting smooth and energy-efficient control:
$c_{u}(u_{h})=u_{h}^{T}Q_{u}u_{h}$ (6)
where $Q_{u}$ is a weighting matrix for control inputs.
@@ V-C Analysis of the Collision Avoidance Cost
$\nabla c_{a}=\sum_{i=1}^{M}\frac{-\beta\lVert v_{h}^{i}\rVert e^{\beta(r_{q}-d_{i})}}{1+e^{\beta(r_{q}-d_{i})}}\cdot\frac{p_{h}-o_{h}^{i}}{d_{i}}$ (7)
The gradient direction points from the obstacle to the waypoint(repulsive direction), and its relationship with the obstacle-directed velocity $v_{h}$ requires specific analysis:
  • If the velocity direction is toward the obstacle ($v_{h}\cdot(p_{h}-o_{h}^{i})<0$), the gradient correction direction is opposite to the velocity direction, thereby mitigating collision;
@@
Furthermore, for distances greater than $r_{q}$, the exponential term $e^{\beta(r_{q}-d_{i})}$ rapidly decays toward zero due to the large choice of $\beta=32$. Consequently, $c_{a}(x_{h})$ is effectively negligible when the obstacle distance exceeds $r_{q}$, This improves optimization efficiency, as the cost is large and decreases quickly within $r_{q}$, while after a few iterations, both the cost and the cost gradient become nearly zero.
Fig. 4 demonstrates how the velocity term in (7) shapes avoidance cost fields. With velocity coupling (Fig. 4a/c), $C_{a}$ creates directional repulsion aligned with motion, pushing waypoints outward from obstacles (gray regions). Without it (Fig. 4b/d), $C_{a}$ forms radial symmetry and traps waypoints in central local minima. Additionally, as shown in the figure, the cost is primarily concentrated within a circular region of radius $r=0.6$ m around the obstacle, ensuring that the cost is negligible beyond this radius.indicating that our cost function is only effective within a specific range of the obstacle.
@@ V-D Quadrotor Dynamics Model
The state transition equation $x_{h+1}=\text{RK4}(f_{d}(\cdot))$ is derived from the quadrotor’s continuous-time dynamics model. In the following subsection, we provide a description of the quadrotor dynamics model used in $f_{d}(\cdot)$.
@@
${\dot{p}}(t)$ $=v(t)$
${\dot{v}}(t)$ $=a(t)-R(t)^{T}DR(t)v(t)$
${\dot{a}}(t)$ $=\frac{K_{a}a_{\text{c}}(t)-a(t)}{\tau_{a}}$
${\dot{\phi}}(t)$ $=\frac{K_{\phi}\phi_{\text{c}}(t)-\phi(t)}{\tau_{\phi}}$ (8)
Here, the matrix $R(t)$ denotes the rotation matrix, computed from the current orientation. The parameter $D=\text{diag}(d_{x},d_{y},d_{z})$ represents linear damping, while $K_{a}$, $\tau_{a}$, $K_{\phi}$, and $\tau_{\phi}$ define the first-order dynamics.
@@ VI-A Implementation Details
To validate the proposed method, we designed both simulation and real-world experiments to evaluate its performance and robustness in unknown and cluttered environments. The setup for the experiments is as follows: the optimizer is limited to a maximum of 10 iterations, and the number of nearest points selected, $M$, is set to 3. The MPC optimization horizon is set to 1 second, with a time step of 0.033 seconds. The repulsion threshold $r_{q}$ was set to 1.0. And the safety distance $d_{s}=0.15$. The NLP problem is solved by CasADi [41] using the IPOPT solver. During each experimental iteration, thirty discrete waypoints are uniformly sampled along the path between the current and goal positions.
@@ VI-B2 AirSim Simulation
We compared our approach with the mapping-and-planning-based Ego-Planner[2] across all scenarios. Quantitative results indicate that while Ego-Planner achieves satisfactory performance at very low speeds ($<3m/s$), its success rate deteriorates markedly with increasing velocity targets - a trend consistent with findings from the original paper’s real-world experiments. Our method maintains consistently high success rates across all tested scenarios and velocity regimes, demonstrating superior robustness and generalization capacity.
@@ VI-C Ablation Experiment
We also conducted experiments to evaluate the impact of depth map and trajectory noise on the performance of our method. We repeated the same experiment in the Flightmare simulator [42], but provided all methods with depth maps computed from stereo pairs and trajectories that were perturbed with Gaussian noise. As shown in the first two columns of Table I, our method is robust to depth map and trajectory noise. The results with noisy inputs are almost identical to those obtained without noise. This demonstrates the robustness of our method to sensor noise and inaccuracies in state estimation.
@@ VI-D Real-World Experiment
[…]
@@ VII Conclusion
Our method demonstrates the effectiveness of directly integrating point clouds into the MPC framework for obstacle avoidance. By leveraging the nearest obstacle points for each waypoint, our approach simplifies the computational pipeline and enhances scalability, enabling high-speed obstacle avoidance in complex environments. The proposed dual KD-tree module further streamlines obstacle detection and path adjustment, ensuring real-time performance and robustness. Our simulations and real-world experiments validate the effectiveness of our approach, highlighting its potential for autonomous navigation in cluttered environments.
Our approach shares similarities in the core structure of the loss function (e.g., penalty terms based on exponential distance decay) of the end-to-end approach [29], and relies on dynamically computing collision loss based on the nearest environmental points. However, our approach explicitly solves the solution through online numerical optimization without the need of training. Although our MPC-based approach still cannot compete this end-to-end approach in agility, our approach stands out as an important alternative which offers better interpretability than the end-to-end approach. Future work will integrate learning-based strategies with optimization-based methods to balance agility, robustness, and interpretability.

### v3

@@ I Introduction
While effective in real-world cluttered environments, this cascade paradigm faces two inherent limitations. First, the accumulated error from each of the sequential modules including occupancy mapping, path planning, trajectory generation, and trajectory tracking could produce biased actions making the quadrotor crash[7]. Second, the sequential workflow also introduces significant computational delays, impairing responsiveness in high-speed scenarios[8]. Both limitations severely restrict the agility and safety of quadrotors. It naturally raises the question of whether approaches without mapping or trajectory generation exist, and whether agility and safety can be significantly enhanced.
@@
Specifically, we introduce a dual KD-tree [15, 16, 17, 18] mechanism, as illustrated in Fig. 1. The Obstacle KD-Tree enables real-time nearest-neighbor queries for collision cost calculation, while the Edge KD-Tree optimizes waypoint adjustment by integrating dilated obstacle boundaries, providing robust initial guesses for the MPC solver to improve optimization efficiency and avoid local minima. This dual-KD tree structure not only simplifies the computational pipeline but also enhances the framework’s ability to handle complex environments, ensuring both dynamic feasibility and real-time performance. Our method eliminates the need for explicit trajectory generation, tracking, or complex mapping processes, significantly enhancing the agility and robustness of autonomous flight in cluttered environments.
@@ III System Overview
The system begins by sparsely sampling a fixed number of waypoints from the current position toward the goal. It then constructs the Obstacle KD-Tree and Edge KD-Tree using the depth map from the RGB-D camera and adjusts the waypoints if a potential collision is detected. The Obstacle KD-Tree stores obstacle points extracted from the depth map. While the Edge KD-Tree is built from a dilated, edge-filtered version of the depth map. Next, the MPC optimization module refines the waypoints generated by the Edge KD-Tree, incorporating kinematic constraints and collision costs. The optimization process minimizes a cost function over a fixed finite time horizon, producing an optimal sequence of control inputs and predictive waypoints. The resulting control commands are then sent to the quadrotor’s low-level flight controller (FCU). This entire process repeats at each time step.
@@ IV Dual KD-Tree Perception Module
Below, we provide a detailed description of the construction and query process of the dual KD-Tree module. The input depth map is first downsampled to a resolution of $64\times 48$ pixels. This resolution was selected as a trade-off between runtime efficiency and sufficient obstacle detail. Two spatially indexed structures are then constructed to enable efficient collision checking and trajectory refinement:
@@ IV-1 Edge KD-Tree and Waypoint Coarse Adjustment
where $\lfloor\cdot\rfloor$ is the floor function, $d_{s}$ is the safe distance (derived from the quadrotor’s radius and depth map noise), $v_{\text{f}}$ is the forward velocity, $\Delta t$ is the MPC optimization time step, and $f_{\text{pix}}$ is the focal length in pixels.
@@ V-A Optimization Formulation
s.t. $\mathbf{x}_{h+1}=\text{RK4}\left(f_{d}\left(\mathbf{x}_{h},\mathbf{u}_{h}\right),\Delta t\right)$
$\mathbf{u}_{\text{min}}\leq\mathbf{u}_{h}\leq\mathbf{u}_{\text{max}}$
$\mathbf{x}_{1}=\mathbf{x}_{\text{init}}$
The discrete-time state vector $x_{h}=[\mathbf{p}_{h},\phi_{h},\mathbf{v}_{h},\mathbf{a}_{h}]^{\top}$ is defined in the inertial frame, where: $\mathbf{p}_{h}\in\mathbb{R}^{3}$ denotes the 3D position, $\phi_{h}\in\mathbb{R}$ is the yaw angle, $\mathbf{v}_{h}\in\mathbb{R}^{3}$ the linear velocity, and $\mathbf{a}_{h}\in\mathbb{R}^{3}$ the linear acceleration. The control input $\mathbf{u}_{h}=[\mathbf{a}_{c,h},\phi_{c,h}]^{\top}$, defined in the inertial frame, comprises $\mathbf{a}_{c,h}\in\mathbb{R}^{3}$ , and $\phi_{c,h}\in\mathbb{R}$ (yaw angle command). The $\mathbf{u}_{\text{min}}$ and $\mathbf{u}_{\text{max}}$ define control input bounds, and $\mathbf{x}_{\text{init}}$ specifies the initial state. The $\text{RK4}(\cdot)$ denotes the fourth-order Runge-Kutta method, while $f_{d}(\cdot)$ represents the quadrotor’s dynamics model. After optimization, only the first action $\mathbf{u}_{1}$ is sent to the low-level flight control unit (FCU) to control the quadrotor.
@@ V-B Cost Function
The cost function $\mathcal{C}\left(\mathbf{x}_{h},\mathbf{u}_{h}\right)$ is a weighted sum of four terms: collision avoidance cost $\mathcal{C}_{a}$, terminal cost $\mathcal{C}_{goal}$, waypoint following cost $\mathcal{C}_{wp}$, and smoothness cost $\mathcal{C}_{u}$. The cost function is defined as:
@@ V-B1 Collision Avoidance Cost
The collision avoidance cost $\mathcal{C}_{a}(\mathbf{x}_{h})$ penalizes the proximity of the h-th waypoint to obstacles. We compute this cost using distances between the quadrotor position $\mathbf{p}_{h}$ and $M$ nearest obstacle points $\{\mathbf{o}_{h}^{i}\}_{i=1}^{M}$ queried from the Obstacle KD-Tree:
$\mathcal{C}_{a}(\mathbf{x}_{h})$ $=\lambda_{c}\sum_{i=1}^{M}\lVert\mathbf{v}_{h}^{i}\rVert\ln\left[1+e^{\beta(r_{q}-\lVert\mathbf{p}_{h}-\mathbf{o}_{h}^{i}\rVert)}\right]$ (3)
@@ V-B2 Terminal Cost
The terminal cost $\mathcal{C}_{goal}\left(\mathbf{x}_{H},\mathbf{u}_{H}\right)$ is used to guide the quadrotor to the final goal point. This cost is computed with the goal state $\mathbf{x}_{\text{goal}}$:
$\mathcal{C}_{goal}\left(\mathbf{x}_{H}\right)=\left(\mathbf{x}_{H}-\mathbf{x}_{\text{goal}}\right)^{T}\mathbf{Q}_{\text{goal}}\left(\mathbf{x}_{H}-\mathbf{x}_{\text{goal}}\right)$ (4)
Here, $\mathbf{Q}_{\text{goal}}$ is a diagonal weighting matrix that adjusts the relative importance of the terminal cost. And the goal state $\mathbf{x}_{\text{goal}}$ is the final waypoint in the sequence.
@@ V-B3 Discrete Waypoint Following Cost
The discrete waypoint following cost, $\mathcal{C}_{wp}(\mathbf{x}_{h})$, encourages the discretized states $\mathbf{x}_{h}$ to closely follow the sparse waypoints $\mathbf{x}_{r}^{h}$:
$\mathcal{C}_{wp}(\mathbf{x}_{h})=\left(\mathbf{x}_{h}-\mathbf{x}_{r}^{h}\right)^{T}\mathbf{Q}_{wp}\left(\mathbf{x}_{h}-\mathbf{x}_{r}^{h}\right)$ (5)
Here, $\mathbf{Q}_{wp}$ is a diagonal weighting matrix. Adjusting $\mathbf{Q}_{wp}$ influences how tightly the quadrotor follows the waypoints.
@@ V-B4 Smoothness Cost
The smoothness cost $\mathcal{C}_{u}(\mathbf{u}_{h})$ penalizes large control inputs $\mathbf{u}_{h}$, promoting smooth and energy-efficient control:
$\mathcal{C}_{u}(\mathbf{u}_{h})=\mathbf{u}_{h}^{T}\mathbf{Q}_{u}\mathbf{u}_{h}$ (6)
where $\mathbf{Q}_{u}$ is a weighting matrix for control inputs.
@@ V-C Analysis of the Collision Avoidance Cost
$\nabla\mathcal{C}_{a}=\sum_{i=1}^{M}\frac{-\beta\lVert\mathbf{v}_{h}^{i}\rVert e^{\beta(r_{q}-d_{i})}}{1+e^{\beta(r_{q}-d_{i})}}\cdot\frac{\mathbf{p}_{h}-\mathbf{o}_{h}^{i}}{d_{i}}$ (7)
The gradient direction points from the obstacle to the waypoint(repulsive direction), and its relationship with the obstacle-directed velocity $\mathbf{v}_{h}\in\mathbb{R}^{3}$ requires specific analysis:
  • If the velocity direction is toward the obstacle ($\mathbf{v}_{h}\cdot(\mathbf{p}_{h}-\mathbf{o}_{h}^{i})<0$), the gradient correction direction is opposite to the velocity direction, thereby mitigating collision;
@@
Furthermore, for distances greater than $r_{q}$, the exponential term $e^{\beta(r_{q}-d_{i})}$ rapidly decays toward zero due to the large choice of $\beta=32$. Consequently, $\mathcal{C}_{a}$ is effectively negligible when the obstacle distance exceeds $r_{q}$, This improves optimization efficiency, as the cost is large and decreases quickly within $r_{q}$, while after a few iterations, both the cost and the cost gradient become nearly zero.
Fig. 4 demonstrates how the velocity term in (7) shapes avoidance cost fields. With velocity coupling (Fig. 4a/c), $\mathcal{C}_{a}$ creates directional repulsion aligned with motion, pushing waypoints outward from obstacles (gray regions). Without it (Fig. 4b/d), $\mathcal{C}_{a}$ forms radial symmetry and traps waypoints in central local minima. Additionally, as shown in the figure, the cost is primarily concentrated within a circular region of radius $r=0.6$ m around the obstacle, ensuring that the cost is negligible beyond this radius.indicating that our cost function is only effective within a specific range of the obstacle.
@@ V-D Quadrotor Dynamics Model
The state transition equation $\mathbf{x}_{h+1}=\text{RK4}(f_{d}(\cdot))$ is derived from the quadrotor’s continuous-time dynamics model. In the following subsection, we provide a description of the quadrotor dynamics model used in $f_{d}(\cdot)$.
@@
$\mathbf{{\dot{p}}}_{h+1}$ $=\mathbf{v}_{h}$
$\mathbf{{\dot{v}}}_{h+1}$ $=\mathbf{a}_{h}-\mathbf{R}_{h}^{T}\mathbf{D}\mathbf{R}_{h}\mathbf{v}_{h}$
$\mathbf{{\dot{a}}}_{h+1}$ $=\frac{\mathbf{K}_{a}\mathbf{a}_{c,h}-\mathbf{a}_{h}}{\mathbf{\tau}_{a}}$
$\mathbf{{\dot{\phi}}}_{h+1}$ $=\frac{k_{\phi}\phi_{c,h}-\phi_{h}}{\tau_{\phi}}$ (8)
Here, the matrix $\mathbf{R}_{h}\in\mathbb{R}^{3\times 3}$ denotes the rotation matrix, computed from the current orientation. The parameter $\mathbf{D}=\text{diag}(d_{x},d_{y},d_{z})$ represents linear damping, while $\mathbf{K}_{a}\in\mathbb{R}^{3\times 3}$, $\mathbf{\tau}_{a}\in\mathbb{R}^{3}$, $k_{\phi}\in\mathbb{R}$, and $\tau_{\phi}\in\mathbb{R}$ define the first-order dynamics, representing the system gains and time constants respectively.
@@ VI-A Implementation Details
To validate the proposed method, we designed both simulation and real-world experiments to evaluate its performance and robustness in unknown and cluttered environments. Experimental parameters are: the optimizer is limited to a maximum of 10 iterations, and the number of nearest points selected, $M$, is set to 3. The MPC optimization horizon is set to 1 second, with a time step of 0.033 seconds. The repulsion threshold $r_{q}=1.0$ is set so that the UAV starts reacting to obstacles when they are within 1.0m, encouraging it to maintain a distance greater than 1.0m from obstacles. This value can be reduced in slower or denser environments. Safety distance $d_{s}=0.15$ was determined according to the physical size of the UAV, which has a radius of approximately 0.1m. The NLP problem is solved by CasADi [40] using the IPOPT solver. Each iteration, thirty discrete waypoints are uniformly sampled between the current and goal positions.
@@ VI-B2 AirSim Simulation
We compared our approach with the mapping-and-planning-based Ego-Planner[2] across all scenarios. Quantitative results indicate that while Ego-Planner achieves satisfactory performance at very low speeds (<3m/s<3m/s), its success rate deteriorates markedly with increasing velocity targets - a trend consistent with findings from the original paper’s real-world experiments. Our method maintains consistently high success rates across all tested scenarios and velocity regimes, demonstrating superior robustness and generalization capacity.
@@ VI-C Ablation Experiment
We also conducted experiments to evaluate the impact of depth map and trajectory noise on the performance of our method. We repeated the same experiment in the Flightmare simulator [41], but provided all methods with depth maps computed from stereo pairs and trajectories that were perturbed with Gaussian noise. The depth maps were generated using the SGM algorithm, which inherently introduces realistic disparity noise, while zero-mean Gaussian noise with a standard deviation of 0.05m was added to trajectory positions to simulate localization errors. As shown in the first two columns of Table I, our method is robust to depth map and trajectory noise. The results with noisy inputs are almost identical to those obtained without noise. This demonstrates the robustness of our method to sensor noise and inaccuracies in state estimation.
@@ VI-D Real-World Experiment
We deployed our method on a lightweight real-world quadrotor platform (weights $0.73\text{\,}\mathrm{kg}\text{/}$) equipped with an Intel RealSense D435i depth camera and an Intel N100 onboard computer (4 cores, 3.4GHz). The platform demonstrates favorable thrust characteristics with a gravity-to-maximum-thrust ratio of 0.35, we integrated a millimeter-level precision motion capture system operating at 200 Hz while the low-level flight control was handled by a Betaflight controller.
@@ VII Conclusion
The proposed method demonstrates the effectiveness of directly integrating point clouds into the MPC framework for obstacle avoidance. By leveraging the nearest obstacle points for each waypoint, this approach simplifies the computational pipeline and enhances scalability, thus enabling high-speed obstacle avoidance in complex environments. The proposed dual KD-tree module further streamlines obstacle detection and path adjustment, ensuring real-time performance and robustness. Simulations and real-world experiments validate the effectiveness of the approach, highlighting its potential for autonomous navigation in cluttered environments.
This approach shares similarities in the core structure of the loss function (e.g., penalty terms based on exponential distance decay) with the end-to-end approach [28], and relies on dynamically computing collision loss based on the nearest environmental points. However, the proposed method explicitly solves the solution through online numerical optimization without the need of training. Although the MPC-based approach still cannot compete this end-to-end approach in agility, it stands out as an important alternative, which offers better interpretability than the end-to-end approach. Future work will integrate learning-based strategies with optimization-based methods to balance agility, robustness, and interpretability.